A 2000 kg truck accelerates from 20kph to 120 kph in 10 seconds. Find its engine’s power output in watts. 


1
Expert's answer
2022-05-11T13:52:46-0400
"P=\frac{W}{t}=\frac{mv^2/2}{t}"

"P=\frac{2000*(120^2-20^2)/3.6^2}{10}=2.16*10^5\:\rm W"